Suppose you held a diversified portfolio consisting of a $7,500 investment in each of 20 different common stocks. The portfolio's beta is 1.73. Now suppose you decided to sell one of the stocks in your portfolio with a beta of 1.0 for $7,500 and use the proceeds to buy another stock with a beta of 1.93. What would your portfolio's new beta be? Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1

Increase in beta as a result of net sales=(1.93-1)=0.93

Hence increase/stock=0.93/20=0.0465

Hence new beta=1.73+0.0465

=1.78(Approx).